### v1

# MaskMoE: Boosting Token-Level Learning via Routing Mask in Mixture-of-Experts

Abstract
Scaling model capacity enhances its capabilities but significantly increases computation. Mixture-of-Experts models (MoEs) address this by allowing model capacity to scale without substantially increasing training or inference costs. Despite their promising results, MoE models encounter several challenges. Primarily, the dispersion of training tokens across multiple experts can lead to underfitting, particularly for infrequent tokens. Additionally, while fixed routing mechanisms can mitigate this issue, they compromise on the diversity of representations. In this paper, we propose MaskMoE, a method designed to enhance token-level learning by employing a routing masking technique within the Mixture-of-Experts model. MaskMoE is capable of maintaining representation diversity while achieving more comprehensive training. Experimental results demonstrate that our method outperforms previous dominant Mixture-of-Experts models in both perplexity (PPL) and downstream tasks.
1 Introduction
Large language models have achieved promising performance on various downstream natural language tasks Touvron et al. (2023a); Dai et al. (2022); Brown et al. (2020); Anil et al. (2023). However, training large language models on extensive textual data has significantly increased computational costs compared to previous works (e.g., BERT Devlin et al. (2019), ELMO Peters et al. (2018), and LSTM Hochreiter and Schmidhuber (1997)). In order to further scale up models within computational budgets, sparse activation networks Child et al. (2019); Du et al. (2022a) receive widespread attention due to their ability to significantly reduce computational costs by using only a few parameters per input. A widely studied approach is Mixture-of-Experts (MoEs)  Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022); Roller et al. (2021), which trains multiple expert networks but selects only a subset to process specific inputs. Compared to dense networks of the same model size, MoEs effectively reduce computational costs.
The current commonly used dynamic routing method in MoEs is to select the expert with the high confidence based on the probability distribution output by an intermediate layer with learnable parameters, i.e., a softmax layer. Previous works Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022) show that MoE models trained based on dynamic routing achieve better performance than dense models with the same amount of training computation.
However, this dynamic routing method introduces a new problem: the routing fluctuation problem Dai et al. (2022). This implies that as training progresses, the same tokens are assigned to different experts in different iterations. Consequently, each expert is trained on only a subset of the same tokens. Additionally, the lack of knowledge sharing and interaction among experts may negatively impact the model’s performance Zhao et al. (2024). Specifically, for the same tokens assigned to different experts, the knowledge each expert learns is harder to share. As previous work Shazeer et al. (2017) has indicated, when there are too many experts, the MoE model exhibits underfitting, and even shows higher perplexity than scenarios with fewer experts. The MoEs models have relatively less impact on the learning of frequent tokens, as these tokens have sufficient data to ensure that each expert receives adequate training. However, the MoEs model disperses infrequent tokens and concepts across various experts, which can lead to underfitting for these tokens. This issue hinders scaling the number of experts for the MoE model. As the number of experts increases while keeping the training data constant, the average number of tokens assigned to each expert further decreases, exacerbating the problem of expert underfitting.
Figure 1: Explanation of random routing masking strategy. For each token, some experts will be randomly masked, with only a subset of experts being visible, e.g., for the token “in” only experts 1, 2 and 3 are visible.
Previously fixed routing strategies Roller et al. (2021) based on Hash Layer, pre-assigning each token to a fixed-route expert, e.g., “deliberate” can only be routed to expert number one as shown in Figure 1. The fixed routing to send the same token to the same expert is more advantageous for the token’s thorough learning, especially for some infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) show that increasing the number of optional experts for the same token input enhances the diversity of representations. Therefore, fixed routing based on the Hash Layer may represent a deficiency in diversity, especially for the representation of frequent tokens. Frequent tokens, in order to achieve diverse representations, require encoding by more experts due to their broader range of usage Koranda et al. (2018). For instance, “the” can follow a variety of nouns, and having a more diverse representation helps to distinguish these nouns, making it easier to predict them in the next token prediction task. To enhance the diversity of representation for frequent tokens, it is necessary to employ more experts to encode them.
In this paper, we propose a routing strategy called MaskMoE. Specifically, we propose a routing masking method to adjust the number of visible experts for tokens of different frequencies by generating a masking vector for each token in the vocabulary before training. For infrequent tokens, the MoE layers employ routing masking to retain only one visible expert to which the token can be routed. For frequent tokens, the MoE layers have a lower routing masking rate, allowing more visible experts. MaskMoE can enhance token-level learning through a routing mask. Specifically, MaskMoE allows the model to learn more intensively about infrequent tokens by routing the same infrequent tokens to the same expert, enabling the expert to be fully trained on those tokens. Meanwhile, MaskMoE maintains the diversity of representations for frequent tokens by still having multiple experts available for routing, which benefits their representation diversity He et al. (2023); Yang et al. (2019). Additionally, despite frequent tokens being routed to multiple experts, their sufficient quantities ensure thorough training. Increase the visible number of experts for high-frequency tokens, i.e., increasing the count of updatable parameters, can further enhance capability. Our experimental results indicate that MaskMoE outperforms previous MoEs models, whether in terms of perplexity or downstream benchmark datasets.
Our contributions are summarized as follows:
  • We propose MaskMoE, which introduces a routing mask method to assign a different number of visible experts to tokens based on their frequency to enhance the model’s token-level learning. MaskMoE ensures sufficient training for infrequent tokens while maintaining diverse representations for frequent tokens.
  • We conduct extensive experiments to analyze the relationship between the number of experts and token frequency. We highlight that dynamic routing, which disperses the same tokens, can lead to underfitting of experts on infrequent tokens, while fixed routing lacks representation diversity for frequent tokens.
  • We validate the effectiveness of the proposed MaskMoE with extensive experiments. Experimental results show that it consistently outperforms previous Mixture-of-Experts models.
2 Related Work
2.1 Language Models
Language models are statistical models designed to optimize the likelihood of token sequences in training data Touvron et al. (2023a). Initially, language models relied on $n$-gram statistics Bahl et al. (1983); Katz (1987); Kneser and Ney (1995). Subsequently, the emphasis shifted to neural network-based models, particularly Recurrent Neural Networks Mikolov et al. (2010) and their variants like LSTMs Graves (2013). These models have demonstrated the ability to learn intricate patterns within textual data, achieving significant success in various language modeling applications.
In recent times, Transformers have become the predominant architecture for language models. Notable examples include BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), GPT-2 Radford et al. (2019), UniLM Dong et al. (2019), and T5 Raffel et al. (2020). The introduction of GPT-3 Brown et al. (2020), which boasts 175 billion parameters, marked a significant milestone due to its exceptional performance across numerous downstream tasks. This has led to a surge in research focusing on large generative language models, including prominent works such as Gopher Rae et al. (2021), PaLM Chowdhery et al. (2022), Pythia Biderman et al. (2023), OPT Zhang et al. (2022), GLM Du et al. (2022b); Zeng et al. (2023) and LLaMA Touvron et al. (2023a, b). Currently, GPT4 OpenAI (2023), trained with Reinforcement Learning from Human Feedback (RLHF), achieves truly remarkable results.
However, as the size of the model grows, the computational demands for both training and inference also increase. MoE models achieve scalability by sparsely activating portions of the model’s parameters, allowing for an increase in model size without significantly raising computational overhead. Consequently, MoE models have been receiving increasing attention recently.
2.2 Mixture-of-Experts
The concept of Mixture-of-Experts (MoE) models is initially proposed by  Jacobs et al. (1991). Later,  Shazeer et al. (2017) applies MoE to LSTM, training an LSTM model with up to 137B parameters. With the rise of the Transformer architecture Vaswani et al. (2017); Devlin et al. (2019), Gshard Lepikhin et al. (2021) first applies MoE to Transformers. Subsequently, powerful MoE models such as GLaM Du et al. (2022a) and Switch Transformer Fedus et al. (2022) emerge.
Early works  Zoph et al. (2022); Fedus et al. (2022); Du et al. (2022a); Lepikhin et al. (2021) mainly focus on dynamic routing methods for learning-to-route MoEs. The Hash Layer  Roller et al. (2021) proposes using a random hashing method to route tokens to a fixed expert, achieving better results than dynamic routing. Then,  Dai et al. 2022 points out that dynamic routing has the problem of routing fluctuation problem, meaning the same input is assigned to different experts as training progresses. Routing fluctuation often harms sample efficiency because the same input updates different experts, which significantly affects infrequent tokens. However, previous works He et al. (2023); Yang et al. (2019) indicate that having more experts allows tokens to obtain richer representations, and fixed routing can affect the diversity of representations, especially for frequent tokens.
Based on the above analysis, we propose MaskMoE, which uses routing masks to alleviate the underfitting problem of infrequent tokens caused by routing fluctuations, while maintaining the representational diversity of frequent tokens.
3 Method
In this section, we first review the training of the MoEs models. Then, we introduce detailed MaskMoE. Finally, we introduce the load balance loss.
3.1 Reviewing the Mixture-of-Experts
The routing of MoEs can be divided into fixed routing Roller et al. (2021); Dai et al. (2022) and dynamic routing with learnable parameters Lepikhin et al. (2021); Du et al. (2022a); Dai et al. (2024); Fedus et al. (2022). The common method for fixed routing is to use random hashing to determine Roller et al. (2021), before training, which experts each token will be sent to. In contrast, dynamic routing employs a routing layer with learnable parameters, i.e., a softmax layer, to decide which expert should process the input tokens.
In terms of structure, the FFN layer of the MoEs model is replaced with sparse expert modules. Given a tokenized input sequence $\mathbf{x}={x_{1},x_{2},...,x_{T}}$ consisting of $T$ tokens, the representation of each token is computed as hidden states by a standard Transformer.
$\mathbf{H}=Tansformer(Embedding(\mathbf{x}))$ (1)
Here, $\mathbf{H}$ denote the hidden states, i.e., $\mathbf{H}=[\mathbf{h}_{1},\mathbf{h}_{2},\ldots,\mathbf{h}_{T}]$. The dense transformer layer consists of self-attention and FFN sublayers.
$\mathbf{\bar{h}}_{t}^{l}=SelfAttn(\mathbf{h_{t}^{l-1}})$ (2)
$\mathbf{h_{t}^{l}}=FFN(\mathbf{\bar{h}}_{t}^{l})$ (3)
Here, $l$ represents the $l$-th transformer layer. MoEs methods replace the FFN sub-layer with the MoEs module. Each MoEs module consists of multiple FFNs, whose outputs are mixed with the gating function $P(\cdot)$, represented by the following formula.
$\mathbf{h_{t}^{l}}=\sum_{i}^{N}P(\mathbf{\bar{h}}_{t}^{l})\cdot FFN_{i}(\mathbf{\bar{h}}_{t}^{l})$ (4)
$N$ represents the number of experts in a single MoE layer. The majority of elements in the distribution of the gated $P(\cdot)$ output are zeros, i.e., only a small portion of experts will be activated. Therefore, increasing the total number of experts in MoEs does not increase computation or inference time.
3.2 Proposed MaskMoE
The typical dynamic routing method routes tokens to the top experts based on the scores output by a softmax layer. This approach to update expert parameters is sparse, meaning that the parameters of unselected experts are not updated. This sparsity may lead to experts underfitting for infrequent tokens. Fixed routing ensures that specific experts consistently process the same input tokens, which helps mitigate the problem of insufficient learning for infrequent tokens. However, it is less effective in fostering a diversity of representations compared to dynamic routing methods. Dynamic routing allows the same token to be encoded by different experts, leading to more varied representations He et al. (2023); Yang et al. (2019). Since frequent tokens can have different meanings depending on the context, fixed routing is less effective than dynamic routing in capturing the nuances of these tokens. To balance the density of parameter updates and the diversity of representations in MoE models, we propose MaskMoE to boost token-level learning via routing mask. As illustrated in Figure 1, MaskMoE achieves this by introducing a routing masking mechanism. For each token, a fixed random masking is applied to the experts, allowing them to be routed only to their respective subsets of experts. The formula is represented as follows.
$\mathbf{p}$ $=\text{Softmax}(W_{g}\mathbf{\bar{h}_{t}^{l}}+\mathbf{M})$ (5)
Where $\mathbf{M}$ is the masking vector used to control the visibility of token-specific experts. For visible experts, the corresponding elements in $\mathbf{M}$ are 0, while for invisible experts, the corresponding elements in $\mathbf{M}$ are $\infty$- ∞. The masking vector is determined before training and does not change during the training process. For multi-layer MoE models, we reuse the same masking vector $\mathbf{M}$ across different MoE layers. Here, $\mathbf{W_{g}}$ is the parameter of the gate. The initial energy masking vector for a token $t$ can be represented by the following formula.
$\mathbf{M}^{t}$ $:=-\infty\cdot\mathbf{1}_{N}$ (6)
$\mathbf{C}$ =$=\{\text{C}_{i}\}_{i=1}^{V}\sim\mathcal{U}(\{0:N-1\})$
$\mathbf{M}_{j}^{t}$ $=0,\quad\forall j\in\mathbf{C}$
Formally, we denote the number of visible experts as $V$. $\mathcal{U}(\{0:N-1\})$ denotes a uniform distribution over the set of integers from 00 to $N-1$. It is worth noting that for infrequent tokens, each token has only one visible expert, i.e., $V=1$. This promotes more thorough learning by the selected expert for such infrequent tokens. For frequent tokens, we also reduce the number of visible experts, but the number of visible experts is greater than 1, i.e., $1<V<=N$. For frequent tokens, it is common for $V$ to be less than $N$. We believe that an appropriately sized $V$ not only allows frequent words to have higher representation diversity compared to fixed routing but also enables more thorough training for the same token compared to dynamic routing.
3.3 Load Balance Loss
MoE models typically require distributed training, where different experts are deployed across various nodes. To avoid load imbalance, where a minority of expert nodes handle the majority of tokens while the majority of expert nodes spend more time idle, this can impact training efficiency. It is generally desirable for the number of tokens processed by different experts to be roughly equal. To achieve load balancing, a load balancing loss is commonly introduced in the training of MoE models. We follow previous works Huang et al. (2024); Fedus et al. (2022) and adopt a widely used loss.
$\mathcal{L}_{bal}$ $=N*\sum_{i=1}^{N}w_{i}\cdot R_{i}$ (7)
[…]
$R_{i}$ $=\frac{1}{T}\sum_{j=1}^{T}\mathbf{p}_{i}^{j}$ (9)
Where $T$ represents the number of tokens in a mini-batch. $\mathbf{p}^{j}$ denotes the probability distribution of the routing output for the j-th token, while $\mathbf{p}_{i}^{j}$ represents the probability value of the i-th expert. It is noteworthy that for infrequent tokens, which have only one visible expert, the load balancing loss does not apply to them. The balancing loss primarily regulates the routing of frequent tokens. Additionally, since infrequent tokens are assigned to experts through random hashing, the overall load is relatively balanced. Therefore, our experiments find that MaskMoE, compared to SMoE, does not experience a significant decrease in training efficiency.
Our final loss is a combination of the language model loss and load-balance loss:
$\mathcal{L}=\mathcal{L}_{lm}+\mathcal{L}_{bal}$ (10)
4 Expertiments
Model Configuration Params Activated Params Pile PPL
Standard Transformer Layers=24, Dense 468M 468M 6.95
SMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.62
Hash Layer Layers=24, MoE_Layers = 1 1.3B 468M 6.56
Share-MoE Layers=24, MoE_Layers = 1 1.3B 468M 6.72
MaskMoE Layers=24, MoE_Layers = 1 1.3B 468M 6.48
SMoE Layers=24, MoE_Layers = 12 10B 468M 6.18
Hash Layer Layers=24, MoE_Layers = 12 10B 468M 6.16
Share-MoE Layers=24, MoE_Layers = 12 10B 468M 6.15
MaskMoE Layers=24, MoE_Layers = 12 10B 468M 6.11
Table 1: Perplexity results of language modeling.
4.1 Pre-training Dataset
Following previous works, we use the Pile dataset Gao et al. (2021) for pre-training data. The Pile dataset is a large-scale, publicly available corpus for language model pre-training, containing 22 domains and over 825GB of English text data. For our experiments, we use the well-known LLaMA tokenizer for tokenization, with a vocabulary size of $32k$. We follow  Xie et al. (2023a); Su et al. (2024) to calculate the sampling rate for each domain based on the number of tokens after tokenization. Due to the computational budget, we follow the pretraining settings of  Xie et al. (2023a); Su et al. (2024); Huang et al. (2024); Xiong et al. (2024); Lian et al. (2024), all models are pre-trained with 100B tokens.
Then, to identify infrequent tokens and frequent tokens, we calculate the frequency of each token in the training set and sort them in descending order of frequency. We categorize the top tokens that cover $P$of the dataset as frequent tokens, and the remaining $(1-P)$as infrequent tokens. Where $P$ is a tunable hyperparameter.
4.2 Experimental Setup
Following  Touvron et al. (2023a, b); Huang et al. (2024), we adopt the LLaMA architecture with 24 transformer layers and a hidden layer dimension of 1024. In line with LLaMA Touvron et al. (2023a), we employ the AdamW  Loshchilov and Hutter (2019) optimizer, and applying a cosine learning rate decay schedule.
For the MoE model, following  Roller et al. (2021), we conduct experiments under both single-layer and multi-layer settings. For the single-layer setting, we replace the final FFN layer with the MoE layer. For the multi-layer MoE models, we follow Gshard Lepikhin et al. (2021), alternating MoE layers every other layer, resulting in a total of $12$ MoE layers in this setting. For single-layer MoE models and dense models, we use a learning rate of $43e-$ 4, while for $12$-layer MoE models, we employ a learning rate of $41e-$ 4 to ensure stable convergence of the models. If not specified otherwise, tokens that account for the top 40% of frequency are considered frequent tokens, while the remaining are considered infrequent tokens, i.e., $P=0.4$. For frequent tokens, the visible expert count $V$ is set to 8, and for infrequent tokens, the visible expert count $V$ is set to 1. All of our implementations are based on the DeepSpeed [^1] https://github.com/microsoft/DeepSpeed library Rajbhandari et al. (2022); Rasley et al. (2020), which offers robust support for MoE-distributed training. By default, we enable the random token selection strategy Kim et al. (2021) implemented in the library to facilitate faster model convergence and better runtime efficiency. Following previous works Roller et al. (2021); Fedus et al. (2022); Dai et al. (2022), unless otherwise specified, our experiments only select the top expert.
4.3 Compared Models
Including MaskMoE, we compare five models for validation experiments. Next, we will briefly introduce five compared models.
  • Dense represents a standard Transformer language model with 468M parameters.
  • SMoE denotes a Switch Transformer Fedus et al. (2022), where the router is a learnable parameter. Unless otherwise specified, each MoE layer has 64 experts.
  • Hash Layer Roller et al. (2021) signifies that through a random hash method, each token is assigned a fixed routing expert prior to training. Unless otherwise specified, each MoE layer has 64 experts.
  • Share-MoE is a hybrid dense and MoE model created using residual connections Rajbhandari et al. (2022). Models with shared experts are currently a popular architecture Dai et al. (2024); DeepSeek-AI et al. (2024); Zhao et al. (2024); Rajbhandari et al. (2022). Unless otherwise specified, Share-MoE has 1 shared expert and 128 routed experts, where each expert is 0.5 times the size of a standard FFN. During both training and inference, in addition to the activation of shared experts, the top expert is also selected from a pool of 128 experts. In such a setup, Share-MoE has the same number of floating-point operations (FLOPs) and activation parameters as SMoE and Hash Layer. This allows for a fair comparison between Share-MoE, SMoE, and Hash Layer.
  • MaskMoE achieves a balance between representation diversity in MoE layers and thorough parameter training through the routing masking method we proposed. Similarly, in this section, MaskMoE also employs the shared expert architecture, which is identical to that of Share-MoE.
Model BoolQ Hellaswag LAMBADA PIQA SIQA StoryCloze Arc-e TriviaQA WebQs
Standard Transformer 56.02 40.73 52.55 67.62 40.79 63.55 51.43 7.44 4.97
MoE_Layers = 1
SMoE 55.26 43.66 53.46 68.28 41.15 64.14 51.98 10.00 6.99
Hash Layer 56.61 43.88 54.67 69.53 41.76 64.19 53.70 9.82 6.89
Share-MoE 58.87 43.00 53.00 68.39 41.45 64.19 51.92 10.54 6.89
MaskMoE 58.38 44.47 55.36 68.99 41.86 65.15 53.14 10.39 7.14
MoE_Layers = 12
SMoE 56.18 46.92 56.37 69.91 41.20 64.99 54.34 13.54 7.33
Hash Layer 57.16 46.61 56.68 71.00 40.83 65.47 55.05 12.97 5.95
Share-MoE 58.71 46.94 56.39 69.69 40.84 65.42 55.89 12.77 6.94
MaskMoE 58.32 47.46 57.46 70.62 41.91 65.69 55.35 15.20 6.74
Table 2: Performances of language models on downstream benchmarks. The best score is marked in bold, and the second best is underlined.
4.4 Main Results
In this section, we first present the model’s perplexity (PPL) on the Pile validation set. Then, following Touvron et al. (2023a); Brown et al. (2020); Su et al. (2024); Dai et al. (2024), we conduct tests on a large number of downstream benchmarks, including zero-shot for BoolQ Clark et al. (2019), HellaSwag Zellers et al. (2019), LAMBADA Paperno et al. (2016), PIQA Bisk et al. (2020), SIQA Sap et al. (2019), StoryCloze Mostafazadeh et al. (2016), and Arc-e Bhakthavatsalam et al. (2021), as well as 5-shot for TriviaQA Joshi et al. (2017) and WebQs Berant et al. (2013). Among them, TriviaQA and WebQs use exact match as the metric, while the remaining benchmarks are evaluated based on accuracy. For a fair comparison, we use the open-source evaluation tool lm-evaluation-harness [^2] https://github.com/EleutherAI/lm-evaluation-harness for assessment.
4.4.1 Perplexity Results
Table 1 shows the main results of language modeling on the Pile validation sets. With the same parameter activations maintained, MoE models consistently achieve improvements compared to the dense model. In comparison to SMoE and Hash Layer, under a single MoE layer setup, MaskMoE has reduced the PPL on the Pile Validation by 0.14 and 0.08 respectively. In a 12-MoE layers setup, the PPL reductions are 0.07 and 0.05 respectively. Compared to Share-MoE, MaskMoE decreases by 0.24 in single MoE layers and by 0.04 in 12-MoE layers. These results demonstrate the rationality of the MaskMoE design.
It is noteworthy that Share-MoE achieves worse results than SMoE and Hash Layer in the single MoE setup. However, MaskMoE consistently maintains superior performance, whether in single or multi-MoE layers. This indicates that MaskMoE exhibits stronger robustness across various settings.
4.4.2 Benchmark Results
As shown in Table 2, we present the performance of the models on downstream tasks. Firstly, we observe that MoE models significantly outperform Dense models in downstream tasks. Then, we find that MaskMoE substantially outperforms the other MoE baselines, whether in single-layer or 12-layers settings.
Compared to SMoE, MaskMoE outperforms in all 9 benchmarks with a single MoE layer and in 8 out of 9 benchmarks with 12-MoE layers. Against Hash Layer, MaskMoE excels in 7 out of 9 benchmarks with a single MoE layer and in 8 out of 9 with 12-MoE layers. Compared to Share-MoE, MaskMoE shows significant improvements in 7 out of 9 benchmarks with a single MoE layer and in 6 out of 9 with 12-MoE layers.
This indicates that the performance enhancement is achieved by using our proposed router mask to control the visibility of experts for different tokens. We believe the improvements are due to two factors. Firstly, compared to SMoE and Share-MoE, infrequent tokens are only routed to a single expert, ensuring more thorough training for these tokens. Secondly, in contrast to the Hash Layer, frequent tokens have multiple expert options, which promotes diversity in the representation of frequent tokens.
5 Analyses
We conduct further experiments to provide more insightful analyses on the proposed MaskMoE. Considering the constraints of computational resources, following previous works Roller et al. (2021); Xie et al. (2023b); Dai et al. (2022), unless otherwise specified, most of the analytical experiments are conducted under the setting of a single MoE layer.
5.1 Impact of Router Mask
$P$ $V_{a}$ $V_{b}$ PPL $P$ $V_{a}$ $V_{b}$ PPL
0.2 8 1 6.518 0.4 16 1 6.509
0.4 8 1 6.506 0.4 4 1 6.511
0.6 8 1 6.518 0.4 64 1 6.549
/ 64 64 6.618$\spadesuit$ / 1 1 6.558 $\clubsuit$
/ 4 4 6.551 / 8 8 6.566
Table 3: Impact of Router Mask. $\spadesuit$ denotes SMoE, $\clubsuit$ denotes Hash Layer. To make the differences in MaskMoE under various configurations more apparent, the values shown in the table are precise to three decimal places.
To visually demonstrate the impact of the routing masking mechanism, the experiments in this subsection do not employ a shared expert structure. Instead, they adopt the same structure as SMoE and Hash Layer, with each MoE layer comprising 64 experts, and the FNN layer is full-sized.
The performance of MaskMoE is predominantly influenced by two hyperparameters: the boundary threshold $P$ for the division of frequent and infrequent tokens, and the maximum number of visible experts $V$ for tokens. As shown in Table 3, we conduct an experimental parameter search for both and reported the Pile Validation PPL. $V_{a}$ represents the number of experts visible for frequent tokens, and $V_{b}$ represents the number of experts visible for infrequent tokens. Notably, when $V_{a}=V_{b}=1$, MaskMoE is equivalent to the Hash Layer, and when $V_{a}=V_{b}=64$, MaskMoE is equivalent to SMoE.
We have the following observations. Firstly, we find that MaskMoE outperforms the Hash Layer. With the $V_{a}=8$ and $V_{b}=1$ settings, MaskMoE reduces perplexity by 0.052 on the Pile validation set compared to the Hash Layer, i.e., $V_{a}=V_{b}=1$. This indicates that MaskMoE’s use of multiple expert encodings for frequent tokens enhances representation diversity, which is beneficial for model performance.
Secondly, compared to SMoE, concentrating tokens into fewer visible experts, such as settings like $V_{a}=V_{b}=8$ or $V_{a}=V_{b}=4$, yields better outcomes. This indicates that routing tokens more densely to a reduced number of visible experts enhances thorough training and improves model performance. Moreover, we observe that the configuration $V_{a}=8$ and $V_{b}=1$ outperforms $V_{a}=V_{b}=8$. This indicates that further reducing the number of visible experts is more beneficial for infrequent tokens.
Lastly, it is noteworthy that for MaskMoE, when $V_{b}=1$, regardless of the value of $V_{a}$ (whether it is 4, 8, 16, or 64), the performance of MaskMoE consistently outperforms that of SMoE and Hash Layer. This demonstrates the robustness of MaskMoE.
5.2 Impact of Shared Expert
To explore the impact of shared experts, in this subsection, we conduct ablation experiments on shared experts, and the results are shown in Table 4. We separately report the average scores of the benchmarks and the PPL of Pile Validation. It is noteworthy that after removing the shared experts, we adopt the same architecture as SMoE and Hash Layer, where the MoE layer consists of 64 fully-sized FFN.
We find that even without shared experts, MaskMoE outperforms SMoE, Hash Layer, and Share-MoE in average benchmark scores and Pile Validation PPL. This further validates the soundness of MaskMoE’s design. Additionally, we observe that incorporating shared expert structures continuously improves MaskMoE’s performance. We believe this is because the shared expert structure promotes greater specialization in token representation DeepSeek-AI et al. (2024).
Model Avg($\uparrow$) PPL($\downarrow$)
SMoE 43.87 6.62
Hash Layer 44.56 6.56
Share-MoE 44.25 6.72
MaskMoE 44.99 6.48
w/o Shared Experts 44.85 6.51
Table 4: Average scores of benchmarks and perplexity on the Pile validation. The best score is marked in bold, and the second best is underlined.
5.3 Number of Experts
(a) Non-Shared Expert Structures
(b) Shared Expert Structures
Figure 2: Comparison of MoE-based Transformers with different numbers of experts.
To investigate the variation in model performance under different configurations of expert numbers, we conducted experiments with 16, 64, and 128 experts. The experimental results, as shown in Figure 2, reveal the following observations.
Firstly, we observe that for SMoE and Share-MoE, increasing the number of experts does not necessarily lead to a lower perplexity. In fact, when the number of experts reaches 128, the model exhibits a higher PPL compared to having only 64 experts. We believe this is because with more experts, the average number of tokens assigned to each expert decreases, leading to further underfitting of the model. In contrast, the performance of the Hash Layer and MaskMoE consistently improves with the increase in the number of experts, indicating that they are more favorable for scaling the number of experts.
Secondly, we observe that compared to the Hash Layer, MaskMoE consistently performs better across all configurations. Although the Hash Layer also enables experts to be adequately trained with the tokens assigned to them, MaskMoE excels in the diversity of representations for frequent tokens. This superiority in handling token diversity contributes to MaskMoE’s overall advantage over Hash Layer.
Lastly, we find that regardless of the number of experts, MaskMoE consistently maintains the best performance. This indicates that MaskMoE has strong parameter robustness.
6 Conclusions
In this paper, we analyze dynamic routing models such as SMoE, which routes the same token to multiple experts, potentially causing the model to underfit, especially for infrequent tokens. While fixed routing can mitigate this issue, it struggles with capturing rich features for frequent tokens. To address these issues, we propose a routing masking method aimed at ensuring adequate learning of infrequent tokens while maintaining rich representations for frequent tokens. Extensive experiments demonstrate that the proposed MaskMoE outperforms various downstream tasks and significantly improves perplexity (PPL) on the Pile dataset compared to SMoE, Hash Layer, and Share-MoE.
7 Limitiations
In this paper, we categorize tokens simply into frequent and infrequent tokens by their frequency. This classification seems somewhat rigid. Considering a smoother classification method could potentially yield better results. Specifically, when tokens are arranged in ascending order, their frequencies demonstrate a gradual decline, and the corresponding number of visible experts also decreases progressively. Due to computational constraints, we do not conduct further experiments. We leave the exploration of smoother partitioning methods for future work.